OpenStudy (anonymous):

First Order Linear Differential Equation Please help with step by step thank you :_)

OpenStudy (anonymous):

\[\frac{ dr }{ d \theta }+ r \tan \theta=\sec \theta \]

OpenStudy (anonymous):

\[r'+r\tan\theta=\sec\theta\] \[r'+r\frac{\sin\theta}{\cos\theta}=\frac{1}{\cos\theta}\] \[r'\cos\theta+r\sin\theta=1\] from inspection and a knowledge that \(\sin^2\theta+cos^2\theta=1\) we can say that \[r = \sin\theta\] is one solution, but also notice that \[r = \sin\theta+A\cos\theta\] is also a solution because \[r'\cos\theta+r\sin\theta = (\cos\theta-A\sin\theta)\cos\theta + (\sin\theta+A\cos\theta)\sin\theta\] \[=\sin^2\theta+cos^2\theta + A\sin\theta\cos\theta-A\cos\theta\sin\theta = \sin^2\theta+cos^2\theta\]

OpenStudy (anonymous):

i solved it with a friend and got the same answer, thanks!